Calculate the pH during an acid base titration when 10.00mL of 0.0111M NaOH is added to a 5.00mL of 0.0222M HF.
Answer #1

- HF is a weak acid. Ko=7.28109 pka = -logka = -log (7.2x104) - ar, pka = 3-14 we know, us= constant v=volume s-concentration